OpenStudy (anonymous):

Use properties of rational numbers to simplify the expression. 21/4 + 32/7 +35/4

OpenStudy (anonymous):

@mathmale @mathmath333 @amistre64 @Compassionate

OpenStudy (amistre64):

to get rid of fractions, we multiply thru by the denominator

OpenStudy (anonymous):

yep and its getting 18.5714285714

OpenStudy (amistre64):

let: 21/4 + 32/7 +35/4 = n 21*4/4 + 32*4/7 +35*4/4 = 4n 21 + 32*4/7 +35 = 4n 21*7 + 32*4*7/7 +35*7 = 4*7n 21*7 + 32*4 +35*7 = 4*7n then just divide off the 4*7 to solve for n

OpenStudy (anonymous):

Wow Thanks You Guys!!

OpenStudy (anonymous):

u r welcome @ MelodyMusic :)

OpenStudy (anonymous):

im still confused :( @shereenkhan @amistre64

OpenStudy (anonymous):

oh wait lol nevermind :)